OpenStudy (anonymous):

lnx=3-ln(x+1) Solve for x.

OpenStudy (anonymous):

ln(x^2-x)=3 x^2-x=e^3 I'm stuck here.

OpenStudy (callisto):

Something is wrong with the sign?

OpenStudy (anonymous):

ln(x^2+x)=3 x^2+x=e^3 x(x+1)=e^3 x=e^3 or e^3-1

OpenStudy (anonymous):

but the answer is 4.01

OpenStudy (callisto):

No... From x^2+x=e^3 The next step is to subtract e^3 from both sides, so you'll get x^2 + x - e^3 =0 Now, use the quadratic formula to solve x.

OpenStudy (anonymous):

why i'm wrong?

OpenStudy (anonymous):

i understand your method. i get 4.01 or 41.2 (rej)

OpenStudy (callisto):

For (x+a)(x+b) = c You can only solve it by putting it as (x+a) = c and (x+b) = c when c=0. The logic is that if you have (x+a)(x+b) = 0, either one of the factor, or both factors, is 0, so you solve it by putting (x+a)=0 or (x+b)=0.

OpenStudy (anonymous):

oh right! thanks a lot! i think i get it.

OpenStudy (callisto):

You're welcome. By the way, may I correct my mistakes? It should be "if you have (x+a)(x+b) = 0, either one of the factor is 0, so you solve it by putting (x+a)=0 or (x+b)=0."